OpenStudy (anonymous):

How do you simplify a radical. For example radical 68

OpenStudy (anonymous):

what try and find two numbers that multiply to be 68, preferably numbers that are perfect squares

OpenStudy (anonymous):

for instance 68 can be written as 4*17

OpenStudy (anonymous):

radical 64 is simplified as 8. radical 25 is simplified as 5. to simply a number that doesnt evenly come out such as 8 or 5, you need to break down your number (68). what is a clean squared number that can go into 68? how about 4? 68 divided by 4 is 17. and 4 can be reduced to radical form of 2. so your answer would be 2radical17

OpenStudy (anonymous):

\[\sqrt{4*17}=\sqrt{4}*\sqrt{17}=2\sqrt{17}\]

OpenStudy (anonymous):

awesome! thank you :)